please show me step by step how to solve 25x^2-16=0
Answer by checkley77(12844)

25x^2-16=0
(5x+4)(5x-6)=0
5x+4=0
5x=-4
x=-4/5 ans.
5x-6=0
5x=6
x=6/5 ans.